# GestureGPT: Toward Zero-shot Free-form Hand Gesture Understanding with Large Language Model Agents

### v4

@@ 7. Conclusion
In this work, we propose GestureGPT, a free-form hand gesture understanding framework that leverages the large language models’ capability to automatically map a spontaneously performed free-form gesture to its targeted interface function. The framework is designed and implemented in the form of triple-agent collaboration, which interprets gesture poses and movements, manages context information, and reason over common sense, to finally discern user intent. In two real-world scenarios, we achieved the highest zero-shot Top-5 gesture grounding accuracy of 83.59% for smart home control and 73.44% for video streaming. The inspiring results showcase, for the first time, the great potential of a free-form hand gesture understanding system, eliminating the need for users to learn, memorize, demonstrate, or manually associate gestures with specific functions. GestureGPT paves way for future research on natural interaction interfaces including but not limited to gestures.
@@ Appendix A Gesture Description Rules
A.1. Training for Threshold and Evaluation for Gesture Description Rules
@@ A.1.1   Training Process
To tune the rule parameters, we first manually annotate the ground truth label for each rule on each gesture class. For each gesture, we assume that most people perform it in the same way, and thus the ground truth label is obtained by common knowledge. (For example, for ‘thumb up’ gesture, we label thumb as straight, index to pinky fingers as bent, index and middle finger are pressed together, etc.)
However, there exist ambiguous cases, in which more than one labels are acceptable. (For example, in ‘peace’ gesture, it is reasonable for thumb to be either straight or bent.) In this case, the label is more than one, and consequently cases like this are not used for parameter tuning.
For most rules are tuned using the whole subsample split. One exception is Finger Flexion on thumb, because in most gesture classes, the ground truth of thumb is either [‘straight’] or [‘straight’, ‘bent’] (thus not used for training), and only one class is [‘bent’]. To address the issue of severe imbalance for thumb flexion, we specially select the training set, which is composed of gesture ‘like’ (thumb is straight) and gesture ‘ok’ (thumb is bent), and each image in this set is manually checked to ensure the quality of the training data).
As is shown above, for those cases used to tune the rules, the ground truth label is one of the two possible states (e.g., ‘straight’ or ‘bent’). But during prediction, we use Three-Way Decision Making method (Yao, 2012). When the rule cannot clearly determine the state, it will predict it as ‘unsure’, to avoid wrong conclusions that may easily mislead LLMs. Consequently, different from the classical binary classification, some modifications have been made to the correctness assessment and tuning criteria to accommodate our specific setting.
@@
  • Unsure: Prediction $=$ ‘unsure’.
  • Error: Prediction $\neq$ ‘unsure’, and prediction $=$ label.
  • Correct: Prediction $\neq$ ‘unsure’, and prediction $\neq$ label.
@@
In ‘unsure’ cases, the loss is between 0 and 1. This is because if we set the loss to be 0 (same as ‘correct’ case), then the rules tend to predict every case to ‘unsure’; if we set the loss to be 1 (same as ‘error’ case), then the rules tend to not predict any ‘unsure’, which may lead to misleading predictions.
A grid search is conducted to find out the optimal parameter with minimal average loss. The optimal parameters for each rule is shown in Table 3
A.1.2   Test Performance
We test the generalization ability of our rules on two dataset: HaGRID test set (third-view) and EgoGesture dataset (first view) (Zhang et al., 2018a).
The performance on HaGRID test set (38576 images) are shown in Table 3. For most rules, the average error rate among all gestures is below 4.2% and the overall accuracy is above 88.2%. One exception is flexion rule for thumb, in which the output ‘unsure’ takes about 45% of all cases. This may be attributed to the unique shape of thumb: the topmost segment of the thumb is curved by nature, so when the thumb is extended, the landmark seems to be slightly bent, which may affect the train and test process. But by Three-Way Decision, we use ‘unsure’ to avoid the misleading information that may potentially confuse the Gesture Description Agent.
@@ A.1.2   Test Performance
To evaluate if the parameters trained on third-view dataset can adapt to first-view images, we tested them on a first-view gesture dataset EgoGesture. We choose 20 gesture classes (fist, measure, zero, one, two, three, four, five, six, seven, eight, nine, ok, three2, C, thumb down, thumb right, thumb left, thumb backward, thumb forward) and label the ground truth for each class. Each gesture has around 250 testing samples. The results are shown in Table 4. (Thumb pointing direction is not evaluated on this dataset because there is no adequate gesture for testing. Only the ‘thumb down’ gesture has a clearly downward pointing direction, yet the images are not pointing strictly downward.)
On this first-view dataset, the error rate of the rules only increase from 2.3% to 6.3% (though the correct rate decrease by around 19% because more ‘unsure’ are predicted), showing that our rules works across different views.
@@
More detailed error analysis of our rules’ performance on EgoGesture Dataset can be found in Table 5. Only those with error rate above 15% is shown.
@@
A.2. Pseudocode for description rules
@@
A.3. Hand State Matrix Details
@@ B.1. Experiment 1: Augmented Reality-Based Smart Home IoT Control
Table 6. Tasks And External Context in the Smart Home Scenario
Instruction External Context
Unlock the Smart Carbinet. [”It is 7:00 PM now.”, ”The child lock on this cabinet supports fingerprint unlocking.”]
Increase the brightness of the light. [”It is 7:05 PM now.”]
Show the next recipes on the smart screen. [”It is 7:12 PM now.”]
Open the oven. [”It is 7:14 PM now.”, ”Recipe instructions: now you need to open the oven”]
Open the air cleaner. [”It is 7:20 PM now.”, ”The air purifier’s sensor detected that the current environment has heavy cooking fumes.”]
Set a timer on the smart screen. [”It is 7:30 PM now.”, ”Recipe instructions: now you need to cook on high heat for five minutes.”]
Switch input source of the smart screen to the smart bell. [”It is 7:32 PM now.”, ”The doorbell is ringing.”]
Make a phone call throught the smart screen. [”It is 7:33 PM now.”, ”Just now, it was the deliveryman delivering goods; the owner of the goods is the user’s roommate, Mark.”]
@@ B.2. Experiment 2: Online Video Streaming on PC
Detailed descriptions of the tasks, including the number of functions and external context information for the video streaming environment, are provided in Table 8. Additionally, the list of functions available for tasks 4, 5, and 6 is presented in Table 9, while the function list for the remaining tasks is shown in Table 10. The variation in the function list is attributed to task 3, which involves full-screening the video page, resulting in a reduced number of available functions.
Figure 11. The experimental platform utilized in the video streaming scenario. (a) and (b) are the video interface and the function distribution, (c) a user was watching the video and do gesture to control.
Table 8. Tasks, Function Numbers, and External Context in the Video Scenario
Instruction Function Numbers External Context
Turn up the volume. 66 [It is 8:01 PM now.]
Drag the progress bar forward. 66 [It is 8:02 PM now. , The user has watched the earlier part of this video.]
Enter full screen mode. 66 [It is 8:04 PM now.]
Pause the video. 17 [It is 8:15 PM now., Right now, the user’s phone is actively receiving an incoming call.]
Resume the video. 17 [It is 8:17 PM now., The user hung up the phone]
Exit full screen mode. 17 [It is 8:48 PM now.]
Like the video. 66 [It is 8:49 PM now.]
Go to the next episode. 66 [It is 8:50 PM now.]
@@ 8 VideoQualitySelection
Table 10. Video Scenario Function List in Task 1,2,3,7,8
ID Name ID Name
0 Redirect to Bilibili Homepage 33 SendButton
1 AnimeSectionLink 34 LikeTheVideo Button
2 LiveBroadcastEntry 35 CoinTipping
3 GameCenterEntry 36 AddToFavorites
4 MemberPurchaseLink 37 Video Share Link Retrieval
5 BilibiliMangaAccess 38 UserProfileLink
6 Match Entry Link 39 GetUserProfile
7 DownloadApp 40 SendMessage
8 SearchButton 41 Follow Button
9 imgOnError 42 AdReportSync
10 OpenVipPage 43 SwitchToggle
11 MessageNotificationLink 44 MarkVideoAsWatched
12 OpenDynamicPage 45 NavigateToEpisode
13 AddToFavorites (first) 46 VideoSelection
14 ViewHistory 47 EpisodeSelection (first)
15 CreativeCenterLink 48 EpisodeNavigation
16 UploadVideoEntry 49 EpisodeSelection (second)
17 VideoProgressControl 50 VideoNavigationLink
18 PlayPauseToggle 51 VideoSeriesNavigation
19 NextButton 52 PlayVideo
20 DisplayPlaybackTime 53 VideoNavigation
21 ChangeVideoQuality 54 OpenVideoLecture
22 ChangePlaybackRate 55 bmgOnLoad (first)
23 SubtitleSettings 56 VideoLink
24 AdjustVolume 57 ProfileLink (first)
25 Video Settings 58 bmgOnLoad (second)
26 PictureInPictureToggle 59 DocumentarySeriesChina
27 ToggleWideScreenMode 60 ProfileLink (second)
28 ToggleWebFullScreen 61 bmgOnLoad (third)
29 ToggleFullscreen 62 VideoPreviewLink
30 DanmakuSwitch 63 DisplayUserProfileLink
31 DanmakuToggleButton 64 VideoLinkNavigation
32 DanmuEtiquetteHint 65 VideoPlayArea

### v5

@@ 7. Conclusion
In this work, we introduce GestureGPT, a human-mimicking framework for understanding free-form hand gestures, which capitalizes on the capabilities of LLMs. Unlike traditional gesture recognition systems that require predefined gestures, GestureGPT automatically maps spontaneously performed, free-form gestures to their targeted interface functions without any user training. This is achieved through a collaborative framework involving three specialized agents: a Gesture Description Agent, a Context Management Agent, and a Gesture Inference Agent. These agents work together to interpret gestures, manage context, and apply common sense reasoning to accurately discern user intents.
Our offline evaluations in two realistic scenarios demonstrated promising results, achieving the highest zero-shot Top-5 gesture grounding accuracy of 83.59% for smart home control and 73.44% for video streaming. These outcomes highlight the significant potential of GestureGPT as a conceptual model for future gesture-based interaction systems. They underscore the viability of our approach in reducing the cognitive load on users by eliminating the necessity to learn, memorize, or manually link gestures to functions.
As a conceptual framework, GestureGPT sets the groundwork for future advancements in natural interaction interfaces. While not yet a practical system ready for everyday use, our framework serves as a foundational step towards realizing more adaptive and intuitive ways to interact with technology. Future research could focus on refining this paradigm, enhancing its responsiveness, and expanding its applicability to include a broader array of natural user interfaces supporting different modalities like facial expressions, body movements, and postures, unleashing the full potential of the framework proposed by GestureGPT.
Acknowledgments
The authors would like to thank Benfeng Xu, whose insightful comments and suggestions were invaluable in writing this paper. This work was supported by the Hunan Provincial Natural Science Foundation of China (No. 2023JJ70009), the HNXJ Philanthropy Foundation (KY24017), the Science and Technology Innovation Program of Hunan Province (No. 2022RC4006), and the Young Scientists Fund of the National Natural Science Foundation of China under Grant No. 62102401.
@@ Appendix A Gesture Description Rules
A.1. Rules Calculation Method
Flexion of a finger is calculated as the total bending angle of each joint. For thumb it is the bending angle of the ip joint, and for other fingers, it is the bending angle of the pip and dip joint. Then, two parameters $straight\_threshold$ and $bent\_threshold$ are set to determine if the finger is straight, bent, or “unsure” if the result falls between them. Since thumb has a different joint structure compared with other fingers, a new pair of thresholds are specially set for thumb.
Proximity of two fingers is calculated as the average minimal distance from each finger’s joint to the other finger. Two thresholds i.e., $together\_threshold$ and $separated\_threshold$ are set to determine if the two fingers are pressed together, separated, or “unsure” if the result falls between them.
Contact of thumb and another finger is computed as the distance between their fingertips. Then, two thresholds, i.e., $contact\_threshold$ and $not\_contact\_threshold$ are set to determine if the two fingers’ fingertips are in contact or not, or “unsure” if the result falls between them.
Point direction of thumb is computed as the direction from thumb’s mcp joint to tip joint. Then, it is compared with two reference vectors representing upward and downward. If a reference vector has the minimal angle with the palm orientation vector and the angle is below $angle\_threshold$, the reference vector would be the thumb’s pointing direction. Note that it is only applicable when thumb is straight. If thumb is bent, the result is set to “unsure”. This is especially useful when discriminating between gestures like “thumb up” and “thumb down”.
Palm orientation is computed as the direction to which the palm is facing. It is computed by the cross product of two vectors within the plane of the palm (Figure 4). Then, the direction is compared with six reference vectors representing upward, downward, left, right, inward and outward. If a reference vector has the minimal angle with the palm orientation vector and the angle is below $angle\_threshold$, the reference vector would be the palm orientation. Otherwise the orientation of palm is set as “unsure”.
A.2. Training for Threshold and Evaluation for Gesture Description Rules
@@ A.2.1. Training Process
To tune the rule parameters, we first manually annotate the ground truth label for each rule on each gesture class. For each gesture, we assume that most people perform it in the same way, and thus the ground truth label is obtained by common knowledge. (For example, for “thumb up” gesture, we label thumb as straight, index to pinky fingers as bent, index and middle finger are pressed together, etc.)
However, there exist ambiguous cases, in which more than one labels are acceptable. (For example, in “peace” gesture, it is reasonable for thumb to be either straight or bent.) In this case, the label is more than one, and consequently cases like this are not used for parameter tuning.
Most rules are tuned using the whole subsample split. One exception is Finger Flexion on thumb, because in most gesture classes, the ground truth of thumb is either [“straight”] or [“straight”, “bent”] (thus not used for training), and only one class is [“bent”]. To address the issue of severe imbalance for thumb flexion, we specially select the training set, which is composed of gesture “like” (thumb is straight) and gesture “ok” (thumb is bent), and each image in this set is manually checked to ensure the quality of the training data.
As is shown above, for those cases used to tune the rules, the ground truth label is one of the two possible states (e.g., “straight” or “bent”). But during prediction, we use Three-Way Decision Making method (Yao, 2012). When the rule cannot clearly determine the state, it will predict it as “unsure”, to avoid wrong conclusions that may easily mislead LLMs. Consequently, different from the classical binary classification, some modifications have been made to the correctness assessment and tuning criteria to accommodate our specific setting.
@@
  • Unsure: Prediction $=$ “unsure”.
  • Error: Prediction $\neq$ “unsure”, and prediction $\neq$ label.
  • Correct: Prediction $\neq$ “unsure”, and prediction = label.
@@
In “unsure” cases, the loss is between 0 and 1. This is because if we set the loss to be 0 (same as “correct” case), then the rules tend to predict every case to “unsure’; if we set the loss to be 1 (same as “error” case), then the rules tend to not predict any “unsure”, which may lead to misleading predictions.
A grid search is conducted to find out the optimal parameter with minimal average loss. The optimal parameters for each rule is shown in Table 3.
A.2.2. Test Performance
We test the generalization ability of our rules on two dataset: HaGRID test set (third-person view) and EgoGesture dataset (first-person view) (Zhang et al., 2018a).
The performance on HaGRID test set (38576 images) is shown in Table 3. For most rules, the average error rate among all gestures is below 4.2% and the overall accuracy is above 88.2%. One exception is flexion rule for thumb, in which the output “unsure” takes about 45% of all cases. This may be attributed to the unique shape of thumb: the topmost segment of the thumb is curved by nature, so when the thumb is extended, the landmark seems to be slightly bent, which may affect the train and test process. But by a three-way decision, we use “unsure” to avoid the misleading information that may potentially confuse the Gesture Description Agent.
@@ A.2.2. Test Performance
To evaluate if the parameters trained on third-person view dataset can adapt to first-person view images, we tested them on a first-person view gesture dataset EgoGesture. We choose 20 gesture classes (fist, measure, zero, one, two, three, four, five, six, seven, eight, nine, ok, three2, C, thumb down, thumb right, thumb left, thumb backward, thumb forward) and label the ground truth for each class. Each gesture has around 250 testing samples. The results are shown in Table 4. (Thumb pointing direction is not evaluated on this dataset because there is no adequate gesture for testing. Only the “thumb down” gesture has a clearly downward pointing direction, yet the images are not pointing strictly downward.)
On this first-person view dataset, the error rate of the rules only increase from 2.3% to 6.3% (though the correct rate decrease by around 19% because more “unsure” are predicted), showing that our rules works across different views.
@@
More detailed error analysis of our rules’ performance on EgoGesture Dataset can be found in Table 5. Only those with error rate above 15% are shown.
@@
A.3. Pseudocode for description rules
@@
A.4. Gesture State Matrix Details
@@ B.1. Experiment 1: Augmented Reality-Based Smart Home IoT Control
Table 6. Task Instructions and External Context in the Smart Home Scenario
Task Instruction External Context
Unlock the Smart Carbinet. [“It is 7:00 PM now.”, “The child lock on this cabinet supports fingerprint unlocking.”]
Increase the brightness of the light. [“It is 7:05 PM now.”]
Show the next recipes on the smart screen. [“It is 7:12 PM now.”]
Open the oven. [“It is 7:14 PM now.”, “Recipe instructions: now you need to open the oven”]
Open the air cleaner. [“It is 7:20 PM now.”, “The air purifier’s sensor detected that the current environment has heavy cooking fumes.”]
Set a timer on the smart screen. [“It is 7:30 PM now.”, “Recipe instructions: now you need to cook on high heat for five minutes.”]
Switch input source of the smart screen to the smart bell. [“It is 7:32 PM now.”, “The doorbell is ringing.”]
Make a phone call throught the smart screen. [“It is 7:33 PM now.”, “Just now, it was the deliveryman delivering goods; the owner of the goods is the user’s roommate, Mark.”]
@@ B.2. Experiment 2: Online Video Streaming on PC
Detailed descriptions of the tasks, including the number of functions and external context information for the video streaming environment, are provided in Table 8. Additionally, the list of functions available for tasks 4, 5, and 6 is presented in Table 9, while the function list for the remaining tasks is shown in Table LABEL:tab:video_function_list_1. The variation in the function list is attributed to task 3, which involves full-screening the video page, resulting in a reduced number of available functions.
Figure 12. The experimental platform utilized in the video streaming scenario. (a) A user was watching the video and performing gesture to control it. (b) and (c) are the video interface and the function distribution.
Table 8. Task Instructions, Function Numbers and External Context in Video Streaming Scenario
Task Instruction Function Numbers External Context
Turn up the volume. 66 [“It is 8:01 PM now.”]
Drag the progress bar forward. 66 [“It is 8:02 PM now.”, “The user has watched the earlier part of this video.”]
Enter full screen mode. 66 [“It is 8:04 PM now.”]
Pause the video. 17 [“It is 8:15 PM now.”, “Right now, the user’s phone is actively receiving an incoming call.”]
Resume the video. 17 [“It is 8:17 PM now.”, “The user hung up the phone”]
Exit full screen mode. 17 [“It is 8:48 PM now.”]
Like the video. 66 [“It is 8:49 PM now.”]
Go to the next episode. 66 [“It is 8:50 PM now.”]
@@ Appendix C System Cost for LLM Use
Appendix D Data Samples and Prompts Utilized in GestureGPT
GestureGPT utilizes a triple-agent collaborative architecture, with each agent guided by its own specific prompt to steer the behavior of the LLM. The gesture data in GestureGPT also undergoes several format transformations. To facilitate better utilization by the community, we will make our prompts and some corresponding data samples publicly available. Due to the length of the system prompt, it has been archived in a GitHub repository. The full prompt can be accessed at https://github.com/studyzx/GestureGPT_ISS.